OpenStudy (anonymous):

Find its vertex its x- and y-intercept(s). (If an answer does not exist, put DNE) f(x)=3x^2+6x-1

OpenStudy (anonymous):

I got the Y-intercept, which is (0,-1).

OpenStudy (anonymous):

Need the smaller and larger values of the X-intercept.

OpenStudy (anonymous):

y-1+3=3(x^2+2x+1) ===> (y+2)=3(x+1)^2 ===> vertex =(-1,-2) ,,, y- intercept is y=1 ,